### v1

# CLIP Meets Diffusion: A Synergistic Approach to Anomaly Detection

Abstract
Anomaly detection is a complex problem due to the ambiguity in defining anomalies, the diversity of anomaly types (e.g., local and global defect), and the scarcity of training data. As such, it necessitates a comprehensive model capable of capturing both low-level and high-level features, even with limited data. To address this, we propose CLIPFUSION, a method that leverages both discriminative and generative foundation models. Specifically, the CLIP-based discriminative model excels at capturing global features, while the diffusion-based generative model effectively captures local details, creating a synergistic and complementary approach. Notably, we introduce a methodology for utilizing cross-attention maps and feature maps extracted from diffusion models specifically for anomaly detection. Experimental results on benchmark datasets (MVTec-AD, VisA) demonstrate that CLIPFUSION consistently outperforms baseline methods, achieving outstanding performance in both anomaly segmentation and classification. We believe that our method underscores the effectiveness of multi-modal and multi-model fusion in tackling the multifaceted challenges of anomaly detection, providing a scalable solution for real-world applications.
1 Introduction
Anomaly detection is an important problem that is widely used in various fields including industry [2, 60] and medicine [47, 35]. In this paper, we address anomaly classification and segmentation.111It is also referred to as anomaly detection and localization, respectively, in the literature. We use the terms anomaly classification and segmentation, and refer to anomaly detection as the overarching problem encompassing both. In anomaly classification (or segmentation), each image (or pixel) is classified as normal or anomalous.
In practice, the number of abnormal images is small, while the potential types of abnormalities are highly diverse, making it challenging to obtain representative examples. Moreover, developing methods that can quickly adapt to new environments (e.g., new processes or products), even with a limited number of normal images, has become a critical research topic [44, 48, 60, 23]. To address this, we focus on the problem within the zero-shot or few-normal-shot222In this paper, ”few-shot” specifically refers to ”few-normal-shot.” setup.
The recent advent of large-scale pretrained foundation models, e.g., GPT [4] and DALL-E [38, 39], has opened new horizons for addressing such challenges. The generalizability and scalability of foundation models make them particularly effective for tackling zero-shot and few-shot tasks such as anomaly detection. In this paper, we present a novel methodology that combines the complementary strengths of discriminative and generative foundation models.
Anomaly detection is a multifaceted problem, and various approaches have been developed based on different perspectives. First, anomaly detection has the aspect of one-class classification. This has led the community to methods that learn normality from normal images and judge what deviates from it as anomalous. Here, many methods extract normal features from normal images using a vision foundation model [10, 11, 43]. Second, anomaly detection has the aspect of two-class classification. The community has employed vision-language foundation models, such as CLIP [37], to compare a query image against normal and abnormal text prompts [25, 29]. Third, anomaly detection also has the aspect of semantic segmentation. Many anomalous features are fine-grained ones that appear in textures or edges. WinCLIP (window-based CLIP) attempted to solve this by sliding windows at multiple scales. However, it is known that CLIP has limitations in learning fine-grained features [51, 26], and such an approach might be suboptimal. Also, it requires hundreds of forward passes due to the exhaustive sliding of multi-scale windows.
We claim that the multifaceted nature of anomaly detection makes it challenging to address using only discriminative models. So, we aim to create synergy by fusing a discriminative vision-language model (VLM) with a generative text-to-image (T2I) model. These are two fundamentally distinct approaches to aligning images and text. Discriminative models, such as CLIP, directly align embeddings in the same space, while generative models, like diffusion models, align them indirectly by jointly performing a common task. Diffusion models learn the relationships between images and texts in a more flexible way, enabling them to capture semantic relationships with greater detail. Therefore, it is essential to leverage both high-level information extracted from the CLIP-based model and low-level information extracted from the diffusion-based model for effective anomaly detection.
We propose CLIPFUSION (= CLIP + Diffusion), a novel method that combines the strengths of multi-modal (language and vision) and multi-model (CLIP and diffusion) approaches. In the zero-shot scenario, CLIPFUSION effectively identifies anomalies by integrating image-text alignment from a CLIP-based model and cross-attention maps [7, 55, 33] from a diffusion-based model. In the few-shot scenario, it further enhances its performance by additionally leveraging feature maps from the CLIP image encoder and the diffusion denoiser, enabling a more comprehensive comparison between the query image and reference images. Figure 1 illustrates qualitative results, showcasing the model’s ability to capture fine-grained anomalies. CLIPFUSION exemplifies a holistic framework that synergizes complementary foundation models, and has the potential to extend beyond anomaly detection and advance other recognition tasks.
Contributions of our work are summarized as follows:
We introduce a novel use of diffusion models as feature extractors for anomaly detection. To our knowledge, this is the first approach to directly construct anomaly maps using cross-attention and feature maps from the diffusion denoiser, making the process fast despite using a diffusion model.
We introduce a novel approach to anomaly detection that jointly leverages CLIP and diffusion models, going beyond previous one-sided approaches. We demonstrate the complementary effect of combining features from discriminative and generative models. Since diffusion models compensate for CLIP’s limitations in capturing fine-grained local details, our method eliminates the need for complex and time-consuming techniques such as sliding windows.
Our method consistently outperforms existing approaches in anomaly detection benchmarks. Unlike most prior works, our method is versatile: it can handle both segmentation and classification tasks under zero-shot and few-shot settings.
Our method is training-free, which makes it robust and easy to apply. In contrast, training-based methods are sensitive to the choice of when to stop training, which is challenging in few-shot setups due to the absence of a validation set.
2 Related Work
Vision(-Language) Models for Anomaly Detection
In works that utilize vision(-language) models in anomaly detection, features are extracted from the intermediate layers of a model pretrained on a discriminative task (e.g., classification or contrastive learning). Anomaly detection is performed by comparing the features of the query image with the features of reference images. Some works utilize ResNet-based models pretrained on ImageNet in the few-shot regime. In SPADE [10], a -nearest neighbors algorithm is employed to identify the reference images closest to the query image, which are then used as the basis for anomaly detection. In PaDiM [11], features extracted from reference images are modeled as Gaussian distributions, and the features of the query image are compared against these distributions. In PatchCore [43], a memory bank is constructed using the features of reference images, and the features of a query image are directly compared against this memory bank. WinCLIP [25] applies local windows and combines the results through harmonic aggregation. PromptAD [29] learns prompt tokens through contrastive training.
A separate line of work assumes access to additional data and leverages it. Some methods require an auxiliary dataset containing both normal and anomaly images, e.g., training on the MVTec-AD dataset and inference on the VisA dataset [22, 59, 52, 32, 5, 58, 16]. AnomalyGPT [17] relies on an auxiliary dataset with a large number of normal samples. In practice, acquiring such data with matching characteristics is challenging, and these methods often struggle when deployed on datasets with distribution shifts. Some methods also rely on ground-truth segmentation masks, which are difficult to obtain in real-world scenarios [22, 5, 58]. MuSc [30] leverages the test set by exhaustively comparing the features of all test images, which is only feasible when the entire test set is available and contains a low proportion of anomalies.
Diffusion Models for Discriminative Tasks
Recently, there has been growing interest in applying diffusion models to discriminative tasks [15]. In DDPM-Seg [1], feature maps are constructed using activations extracted from intermediate layers of the denoiser and are utilized for semantic segmentation. VPD [56] further incorporates cross-attention maps for semantic segmentation and depth estimation. These methods require fine-tuning for downstream tasks. On the other hand, DiffusionClassifier [27] performs zero-shot classification by determining which text conditioning best predicts the noise. P2PCAC [20] and USCSD [19] reveal that cross-attention maps contain rich information about spatial layouts through image editing and semantic correspondence tasks.
3 Preliminaries
CLIP
CLIP [37] consists of an image encoder and a text encoder . These encoders map images and texts into the same embedding space. The model is trained to match correct image-text pairs with high similarity while assigning low similarity to incorrect pairs. This process is optimized using the following contrastive loss [8]:
where is the set of candidate text prompts, is the temperature hyperparameter, and is the cosine similarity measure.
Diffusion Models
Diffusion models [21] are generative models that produce data by reversing a gradual noising process. In the forward process, Gaussian noise is incrementally added to clean data over multiple steps, transforming it into pure noise. The reverse process is learned via a neural network that denoises the data step-by-step to recover the original signal. When a text prompt is provided, it acts as a guide so that the image being generated matches the meaning or description in the text. They can also be utilized for inpainting a specific region of an image. At each step, the untouched region is preserved by replacing it with the corresponding part of the original image, ensuring that only the specified region is generated and seamlessly blended with the rest of the image [40]. See Appendix A for more details.
4 Method
The final model, CLIPFUSION, is a fusion of our simple CLIP-based model and our complementary diffusion-based model. For convenience, we call the CLIP-based model PatchCLIP (patch-based CLIP) and the diffusion-based model MapDiff (map-based diffusion). We first describe the characteristics of each model and what resources can be extracted from them for later use.
PatchCLIP
CLIP processes an image by dividing it into multiple patches and outputs both patch embeddings and a class token embedding. While the class token embedding is trained to be aligned with text embeddings, we observe that patch embeddings are also beneficial, so we leverage both. Given that fine-grained details are handled by the diffusion-based model, PatchCLIP provides a simple yet effective spatial prior, avoiding the cumbersome multi-scale window sliding in WinCLIP.
Additionally, we utilize features extracted from CLIP’s image encoder. To achieve this, we make use of the VV-attention mechanism proposed in CLIP surgery [31, 29]. This mechanism was introduced to complement CLIP’s QK-attention mechanism, which is specialized in global feature extraction. It facilitates local feature extraction by enabling direct interactions between values (hence the name VV-attention) that carry information about the patches themselves.
MapDiff
We extract and utilize cross-attention maps and feature maps from the diffusion denoiser. To achieve this, we use an inpainting pipeline that takes a clean image as input, rather than a generation pipeline that starts with random noise. In the inpainting pipeline, a mask is provided to specify the region where noise will be added. Here, we use the trick of applying an empty mask. This prevents noise from being added, and the denoiser perceives the original image as noisy, allowing us to control the timestep to extract features at the desired level.
By passing the clean image through the denoiser once, we extract the maps from its intermediate layers. A large timestep makes the denoiser interpret the image as highly noisy, capturing high-level pattern information, while a small timestep emphasizes finer details.
4.1 Anomaly Segmentation
The final anomaly score map for segmentation is obtained by fusing the score maps from PatchCLIP and MapDiff. In the zero-shot setup, only the language-guided score maps and are used. In the few-shot setup, the vision-guided score maps and are additionally incorporated. The final anomaly score map is computed as follows:
where is the weight between the models. In the zero-shot setup, .
4.1.1 Zero-shot
In the zero-shot setup, given a query image, we leverage the relationships between images and texts learned by multi-modal foundation models to create anomaly score maps.
CLIP Language-guided Score Map
The query image is input to the CLIP image encoder (e.g., ViT), which divides it into patches, processes them, and outputs patch embeddings. On the other hand, normal and abnormal text prompts are input to the CLIP text encoder and normal and abnormal text embeddings are output. Then, the degree (Equation 1) to which each patch embedding is aligned to the abnormal text embedding is calculated for all patches and organized into a map. This is upsampled to the original image size through interpolation to obtain .
Diff Language-guided Score Map
The query image is passed to the diffusion inpainting pipeline along with an empty mask, a timestep, and a text prompt. The text prompt includes an object word and a state word, for example:
“a photo of a [object] with [state]."
where [object] refers to the entity (e.g., “bottle") in the image, and [state] describes a possible abnormal condition (e.g., “crack") of it. The [state] token interacts with the image as it passes through the denoiser (e.g., U-Net), generating a cross-attention map. A higher value at a specific location in the map indicates that the location responds more strongly to the [state] token (i.e., the abnormal condition). Therefore, we can use this cross-attention map as an anomaly score map (Figure 3). Additionally, we calculate the average of cross-attention maps derived using various possible abnormal conditions (e.g., “crack", “hole", “residue", “damage") to obtain .
This can be interpreted as follows. In anomaly segmentation, given an image , we are interested in the probability that the -th pixel is abnormal (). This probability can be expressed as follows by conditioning on a text prompt :
where represents the set of text prompts that describe abnormal conditions. Assuming that the text prompt is uniform and interpreting as the cross-attention map intensity at the corresponding location, we have:
4.1.2 Few-shot
In the few-shot setup, normal reference images are utilized. For the -shot case, images are used to construct feature memory banks, which are then compared with the features of the query image.
CLIP Vision-guided Score Map
The reference images are passed through the CLIP’s image encoder to extract reference features and stored in the memory bank. Features with low-level semantic information are extracted from a middle encoder block, while those with high-level semantic information come from a later encoder block.
The set of reference features extracted from the encoder block is denoted as . For a given encoder block , the feature map extracted from the query image is denoted as . Then, is calculated by comparing the feature with the reference features , as follows:
where is the set of blocks . That is, a score is computed based on the similarity, and the smallest score among all reference features in is conservatively used as the anomaly score for . Finally, the average over is taken as the final anomaly score.
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPRO | AUROC | AUPRO | ||
| 0-shot | Trans-MM [6] | 57.50.0 | 21.90.0 | 49.40.0 | 10.20.0 |
| MaskCLIP [57] | 63.70.0 | 40.50.0 | 60.90.0 | 27.30.0 | |
| WinCLIP [25] | 85.10.0 | 64.60.0 | 79.60.0 | 56.80.0 | |
| CLIPFUSION (ours) | 90.90.0 | 87.10.0 | 92.10.0 | 83.30.0 | |
| 1-shot | PaDiM [11] | 89.30.9 | 73.32.0 | 89.90.8 | 64.32.4 |
| PatchCore [43] | 92.01.0 | 79.72.0 | 95.40.6 | 80.52.5 | |
| WinCLIP [25] | 95.20.5 | 87.11.2 | 96.40.4 | 85.12.1 | |
| PromptAD [29] | 94.90.8 | 87.71.5 | 95.70.9 | 84.01.2 | |
| CLIPFUSION (ours) | 95.80.1 | 90.40.1 | 97.30.1 | 87.00.3 | |
| 2-shot | PaDiM [11] | 91.30.7 | 78.21.8 | 92.00.7 | 70.12.6 |
| PatchCore [43] | 93.30.6 | 82.31.3 | 96.10.5 | 82.62.3 | |
| WinCLIP [25] | 96.00.3 | 88.40.9 | 96.80.3 | 86.21.4 | |
| PromptAD [29] | 95.00.7 | 88.31.0 | 96.30.9 | 84.80.8 | |
| CLIPFUSION (ours) | 96.30.1 | 91.40.2 | 97.70.1 | 87.30.2 | |
| 4-shot | PaDiM [11] | 92.60.7 | 81.31.9 | 93.20.5 | 72.61.9 |
| PatchCore [43] | 94.30.5 | 84.31.6 | 96.80.3 | 84.91.4 | |
| WinCLIP [25] | 96.20.3 | 89.00.8 | 97.20.2 | 87.60.9 | |
| PromptAD [29] | 95.50.6 | 89.00.6 | 96.30.8 | 85.10.9 | |
| CLIPFUSION (ours) | 96.80.1 | 92.10.2 | 98.00.1 | 87.70.5 | |
Diff Vision-guided Score Map
Reference images are passed through the diffusion denoiser to extract features, which are stored in the memory bank. The denoiser consists of three parts: an encoder (down-blocks), a bottleneck (mid-block), and a decoder (up-blocks). In the encoder, during compression, information is passed to the decoder via skip connections. The bottleneck is the most compressed stage, where much of the detailed information is lost, making it unsuitable for feature extraction. Therefore, we extract features from the decoder. Specifically, we focus on the middle and later decoder blocks, as they provide the most semantically rich features [1, 56].
When timestep is given, the set of reference features extracted from the decoder block is denoted as . The feature map extracted from the query image at timestep and decoder block is denoted as . Then, is calculated as follows:
where is the set of combinations .
4.2 Anomaly Classification
The final anomaly score for classification is obtained by fusing four scores:
In the zero-shot setup, only the language-guided scores are used. Specifically, the CLIP language-guided score is defined as the degree (Equation 1) of alignment between the query image’s class token embedding and the abnormal text embedding. The diffusion language-guided score aggregates the diffusion language-guided score map and is defined as follows:
This means that the score becomes larger when the degree of focus of the cross-attention map is high, that is, when the maximum value is significantly greater than the median value.
In the few-shot setup, we additionally aggregate the vision-guided score maps, defining the scores as follows in accordance with common practice [25]:
5 Comparisons
Datasets
We use the MVTec-AD [2] and VisA [60] datasets, which are mainly used as benchmark datasets in anomaly detection. Each dataset consists of several object categories, and each object category is divided into a training set containing only normal images and a test set containing a mixture of normal and abnormal images. The abnormal images have various types of anomalies. Query images are from the test set, and reference images are sampled from the training set.
Baseline methods
We compare against methods with the same setup, that is, methods that use only the provided zero-shot or few-shot normal images and do not rely on any additional supervision. This setup reflects a realistic and challenging scenario commonly encountered in practice.
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPR | AUROC | AUPR | ||
| 0-shot | CLIP-AC [37] | 74.00.0 | 89.10.0 | 59.30.0 | 67.00.0 |
| WinCLIP [25] | 91.80.0 | 96.50.0 | 78.10.0 | 81.20.0 | |
| CLIPFUSION (ours) | 93.60.0 | 97.00.0 | 79.50.0 | 83.60.0 | |
| 1-shot | PaDiM [11] | 76.63.1 | 88.11.7 | 62.85.4 | 68.34.0 |
| PatchCore [43] | 83.43.0 | 92.21.5 | 79.92.9 | 82.82.3 | |
| WinCLIP [25] | 93.12.0 | 96.50.9 | 83.84.0 | 85.14.0 | |
| PromptAD [29] | 92.61.0 | 95.70.9 | 83.12.1 | 85.71.2 | |
| CLIPFUSION (ours) | 95.40.8 | 97.80.4 | 86.50.1 | 88.40.2 | |
| 2-shot | PaDiM [11] | 78.93.1 | 89.31.7 | 67.45.1 | 71.63.8 |
| PatchCore [43] | 86.33.3 | 93.81.7 | 81.64.0 | 84.83.2 | |
| WinCLIP [25] | 94.41.3 | 97.00.7 | 84.62.4 | 85.82.7 | |
| PromptAD [29] | 93.21.0 | 96.11.0 | 85.01.8 | 87.21.8 | |
| CLIPFUSION (ours) | 96.30.3 | 98.20.2 | 87.10.5 | 88.80.5 | |
| 4-shot | PaDiM [11] | 80.42.5 | 90.51.6 | 72.82.9 | 75.62.2 |
| PatchCore [43] | 88.82.6 | 94.51.5 | 85.32.1 | 87.52.1 | |
| WinCLIP [25] | 95.21.3 | 97.30.6 | 87.31.8 | 88.81.8 | |
| PromptAD [29] | 93.80.7 | 96.50.7 | 86.21.6 | 88.11.9 | |
| CLIPFUSION (ours) | 96.90.2 | 98.60.1 | 88.10.2 | 89.70.1 | |
Specifically, we compare our method, CLIPFUSION, with methods utilizing vision models (PaDiM, PatchCore) and those utilizing vision-language models (CLIP-AC, Trans-MM, MaskCLIP, WinCLIP, PromptAD333The authors of PromptAD use the test set as a validation set during training to record the best performance for each category separately and report the aggregated results. For a fair comparison, we treat the training epoch as a fixed hyperparameter across categories and report the best-performing checkpoint reproduced using the official codebase. See Appendix F for more details.). Note that, among these methods, only CLIPFUSION and WinCLIP are fully capable of handling both segmentation and classification tasks in both zero-shot and few-shot scenarios.
Experimental details
5.1 Anomaly Segmentation
Table 1 presents a comparison of our method’s performance with baseline methods under zero-shot and few-shot setups for anomaly segmentation. The results show that CLIPFUSION consistently outperforms the baseline methods across all setups, datasets, and metrics. Our method, in particular, demonstrates dramatic performance improvements in the zero-shot setup. This highlights the critical contribution of the diffusion-based model’s cross-attention map to zero-shot anomaly segmentation. Additionally, the significant performance improvements in the few-shot setup demonstrate how the diffusion-based model complements the CLIP-based model.
5.2 Anomaly Classification
Table 13 shows the performance comparison between our method and baseline methods for anomaly classification. CLIPFUSION consistently outperforms all baseline methods. This indicates that the diffusion-based model enhances performance not only in the segmentation task but also in the classification task by reinforcing local information.
In both segmentation and classification tasks, the consistent performance improvements as the number of shots increase highlight the scalability of our method with minimal additional data. Furthermore, the smaller standard deviations observed in our results indicate the robustness of our approach. This robustness can be attributed to the integration of multiple models, which mitigates the variability.
6 Empirical Study
In this section, all results are reported under the one-shot setup for consistency.
| Task | Model | Dataset | ||
| CLIP | Diff | MVTec-AD | VisA | |
| Seg. | ✓ | × | 94.7 | 93.7 |
| × | ✓ | 95.1 | 96.5 | |
| ✓ | ✓ | 95.8 | 97.3 | |
| Cls. | ✓ | × | 93.5 | 82.2 |
| × | ✓ | 87.2 | 78.9 | |
| ✓ | ✓ | 95.4 | 86.5 | |
| Task | Modality | Dataset | ||
| Language | Vision | MVTec-AD | VisA | |
| Seg. | ✓ | × | 90.9 | 92.1 |
| × | ✓ | 95.4 | 96.1 | |
| ✓ | ✓ | 95.8 | 97.3 | |
| Cls. | ✓ | × | 93.2 | 79.5 |
| × | ✓ | 90.5 | 82.8 | |
| ✓ | ✓ | 95.4 | 86.5 | |
| Task | Configuration | MVTec-AD | VisA |
| Seg. | Step , Block | 94.3 | 95.1 |
| Step , Block | 95.1 | 96.5 | |
| Cls. | Step , Block | 86.3 | 76.2 |
| Step , Block | 87.2 | 78.9 |
| Task | Early Block | MVTec-AD | VisA |
| Seg. | w/ | 94.9 | 93.7 |
| w/o | 95.1 | 96.5 | |
| Cls. | w/ | 86.5 | 74.9 |
| w/o | 87.2 | 78.9 |
6.1 Ablation Study
Table 3 presents the comparative performance of different models. As expected, the diffusion-based model, which captures local features effectively, performs better in the segmentation task, while the CLIP-based model, which captures global features effectively, performs better in the classification task. The combination of different models consistently yields better results compared to using either model alone. This empirical evidence suggests that the two models capture complementary features, making their integration highly beneficial for both segmentation and classification tasks.
Table 4 presents the comparative performance of different modalities. The results highlight the complementary strengths of both modalities. Combining the two modalities enhance performance across tasks by leveraging both semantic and spatial information.
6.2 Extraction Strategies from Diffusion Denoisers
We explore strategies to effectively extract features from diffusion denoisers. Here, we conduct experiments using only MapDiff to see the effect more clearly.
Table 5 highlights the impact of matching timestep and block numbers when calculating (Equation 6). Matching higher timesteps with lower blocks yields better performance, and vice versa (Timestep 201 → Block 3, Timestep 401 → Block 2, Timestep 801 → Block 1). This pattern reflects the intrinsic behavior of the diffusion denoiser: at higher timesteps, the denoiser interprets the input image as highly noisy and prioritizes the extraction of global features. The lower blocks, which process coarse-grained information, are particularly well-suited for this task. Using higher timesteps can be seen as a regularization strategy, as they suppress local details and highlight broader patterns.
Table 6 shows that when calculating (Equation 6), including the early block (Block 0) does not enhance the model’s performance. On the contrary, excluding this block consistently yields better results. This suggests that the early stages of the decoder act as preparatory steps for the later stages, without forming concrete features.
6.3 One-for-all paradigm
| Method | AUROC | AUPR |
| IIPAD | 94.2 | 97.2 |
| CLIPFUSION | 94.5 | 97.3 |
IIPAD [34] recently proposes a new setting for few-shot anomaly detection, termed the one-for-all paradigm. It aims to train a single unified model that can detect anomalies across multiple object categories. In this paradigm, the one-shot setting refers to using normal images (one for each category), where is the number of categories. Table 7 presents a comparison with IIPAD (the SOTA method under this paradigm) on the MVTec-AD dataset in the one-shot setting, where our method achieves higher performance. See Appendix G for more results.
6.4 Inference speed
| Latency [ms] | 0-shot | 1-shot | 2-shot | 4-shot |
| WinCLIP | 4782 | 9514 | 9633 | 9710 |
| CLIPFUSION | 783 | 1240 | 1269 | 1278 |
Our method can be viewed as an extension of WinCLIP. We report the inference latency measured on an NVIDIA A100 GPU in Table 8, with a batch size of 1. As shown, our method is approximately 6–8 times faster than WinCLIP. To capture fine-grained details, WinCLIP slides multi-scale windows over the query image, resulting in hundreds of forward passes. In contrast, our CLIP-based model utilizes patch embeddings obtained from a single forward pass. On top of this, we employ a diffusion-based model not as a denoiser for iterative generation, but as a feature extractor to capture fine-grained details. As a result, our method requires only a few forward passes in total.
7 Conclusion
We propose a novel methodology that synergistically combines discriminative and generative foundation models, leveraging their complementary strengths. As a concrete application, we introduce CLIPFUSION, which integrates CLIP and diffusion models for anomaly detection and consistently outperforms baselines in segmentation and classification across zero-shot and few-shot scenarios. The holistic nature of this approach not only advances anomaly detection but also demonstrates the potential of multi-modal and multi-model integration for complex recognition tasks.
Appendix A Further preliminaries on diffusion models
A diffusion model [21] is trained through a sequential process that involves forward diffusion and backward denoising. In the diffusion process, random noise is progressively added to a clean image over several steps, following the equation:
where is the noise schedule hyperparameter, which determines the amount of noise added to the image [36]. In the denoising process, random noise is progressively removed over several steps using the following equation to restore the original image:
In this process, the denoiser learns to predict the added noise from the noisy image at a given time step [49]. During inference, random noise is sampled, and the denoiser iteratively removes it to generate a new image. By conditioning on a text prompt , the denoiser guides the generation process to align the resulting image with the provided text. Diffusion models demonstrate superior performance in understanding image details by generating high-quality images compared to other generative models [13].
Appendix B Further motivation for feature-based anomaly detection
Addressing anomaly detection with a single-perspective approach is challenging due to the open-ended nature of abnormality. Discriminative models focus on global features but often miss fine-grained details, which are crucial for detecting local structural or textural anomalies.
When utilizing generative models in anomaly detection, the reconstruction-based approach has traditionally been the dominant paradigm [53, 54, 18]. This approach compares the original image and the reconstructed image at both ends. However, since generative models are primarily designed for creative image generation, they can produce high reconstruction errors even for normal images. In addition, deciding how to apply the mask, including its location and size is challenging.444Since the location of the anomalous part is not known, it is unclear where to apply the mask. Also, if the mask is too large, the reconstructed image may deviate significantly from the original image. Conversely, if the mask is too small and fails to fully cover the anomaly, the anomaly may persist or even extend in the reconstructed image. As a result, the feature-based approach has recently gained prominence.
Building on this shift, we propose that a more fundamental approach to utilizing generative models for anomaly detection is to extract features midway through the process. This leverages the image understanding capability that underlies their image generation capability.
Appendix C Further experimental details
Data pre-processing
For PatchCLIP, we use the same pre-processing pipeline as in WinCLIP [25], which is specified in OpenCLIP [24], for both MVTec-AD and VisA datasets. First, bicubic interpolation is used to resize images to a resolution of 240 to fit the ViT-B-16-plus-240 model. After which, channel-wise standardization is applied with the precomputed values of (0.48145466, 0.4578275, 0.40821073) for the mean and (0.26862954, 0.26130258, 0.27577711) for the standard deviation. For MapDiff, we use bilinear interpolation to resize images to a resolution of 512 before feeding into a customized Stable Diffusion Inpainting pipeline from the diffusers library [50].
Prompts
The prompts are derived from those used to train CLIP for the ImageNet dataset [12]. For example, in the case of a segmentation task on MVTec-AD, PatchCLIP uses the prompt a good, cropped picture of the [state] [object] for classification," where [state] is either “damaged" or “perfect." MapDiff uses the prompt a close-up cropped png photo of a [object] with [state] for anomaly segmentation," where [state] includes “crack," “hole," “residue," and “damage," for query images, and “a photo of a perfect [object]" for reference images.
Evaluation metrics
We follow previous studies [10, 11, 43, 25] and use the following performance metrics. For anomaly classification, we report the Area Under the ROC Curve (AUROC) at the image level and the Area Under the Precision-Recall Curve (AUPR). For anomaly segmentation, we report the AUROC at the pixel level and the Area Under the Per-Region Overlap Curve (AUPRO) [3, 28]. Both AUPR and AUPRO are particularly sensitive to class imbalance.
Baseline performance
The performance of Trans-MM, MaskCLIP, PaDiM, PatchCore, CLIP-AC, and WinCLIP are referenced from the WinCLIP paper [25].
Appendix D Further empirical study
Table 9 demonstrates the effectiveness of extracting cross-attention maps from the encoder (down-blocks) and decoder (up-blocks) while excluding the bottleneck (mid-block). Integrating the bottleneck results in suboptimal performance due to its lack of spatial granularity critical for tasks like anomaly detection.
| Task | Denoiser Component | Dataset | |||
| Enc | Bneck | Dec | MVTec-AD | VisA | |
| Seg. | × | × | ✓ | 94.4 | 96.0 |
| ✓ | ✓ | ✓ | 94.4 | 96.4 | |
| ✓ | × | ✓ | 95.1 | 96.5 | |
| Cls. | × | × | ✓ | 83.2 | 74.1 |
| ✓ | ✓ | ✓ | 86.8 | 76.1 | |
| ✓ | × | ✓ | 87.2 | 78.9 | |
Table 10 illustrates the impact of using state ensemble for a cross-attention map. Without ensemble, only a generic term “damage" is used to generate the cross-attention map. In contrast, the ensemble approach averages cross-attention maps derived from both the generic term “damage" and more specific terms such as “crack," “hole," and “residue." This approach significantly improves performance across tasks. By incorporating these multiple specific states, the method generates diverse cross-attention maps that capture fine-grained and complementary aspects of anomalies.
| Task | Ensemble | Dataset | |
| MVTec-AD | VisA | ||
| Seg. | w/o | 94.0 | 95.8 |
| w/ | 95.1 | 96.5 | |
| Cls. | w/o | 85.4 | 73.1 |
| w/ | 87.2 | 78.9 | |
Table 11 shows the impact of customizing state descriptors for each object class on performance. For example, for both Cable and Pill, replacing the default set of states with a class-specific set composed of frequently occurring anomaly types leads to consistent performance improvements.
| Object | State Set | AUROC |
| Cable | “crack", “hole", “residue", “damage" | 94.6 |
| “crack", “poke", “scratch" | 95.0 | |
| Pill | “crack", “hole", “residue", “damage" | 83.0 |
| “crack", “scratch", “residue" | 83.5 |
Appendix E Limitations
Our method leverages pretrained foundation models (e.g., CLIP and Stable Diffusion) without any task-specific fine-tuning. While this reduces adaptability to domain-specific anomalies, the method still generalizes well in zero-shot and few-shot scenarios across diverse object categories. Investigating how fine-tuning these foundation models for specific use cases could further improve performance can be an interesting direction for future research.
Appendix F Limitations of few-shot training
Training-based methods inherently suffer from epoch selection sensitivity. Figure 4 shows that the epoch at which PromptAD performs best differs significantly across object categories. In conventional training setups, such epoch selection is guided by a validation set. However, in few-shot anomaly detection, where only one to four images are available, it is practically hard to reserve any for a separate validation set. The authors of PromptAD circumvent this by using the test set for validation in their official implementation.555https://github.com/FuNz-0/PromptAD However, test sets are not accessible during model training in real-world applications. In contrast, CLIPFUSION is training-free and uses few-shot examples solely for feature extraction, making validation sets and epoch selection unnecessary.
Figure 5 reveals that even when the training epoch is fixed, performance varies considerably across different random seeds. This highlights the instability and seed sensitivity of few-shot training, making it difficult to ensure consistent and reproducible results without a validation set.
Appendix G Further results in the one-for-all paradigm
We evaluate anomaly detection performance under the one-for-all paradigm, where a single model is applied across all object categories. Recall that, in this paradigm, the one-shot setting provides one image per category, for a total of images across object categories. IIPAD trains a prompt generator using these examples for the one-for-all setting. In contrast, CLIPFUSION remains entirely training-free. Nevertheless, it outperforms or matches IIPAD across all configurations and datasets for both segmentation and classification.
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPR | AUROC | AUPR | ||
| 1-shot | IIPAD | 96.4 | 89.8 | 96.9 | 87.3 |
| CLIPFUSION (Ours) | 96.0 | 91.1 | 97.2 | 87.3 | |
| 2-shot | IIPAD | 96.7 | 90.3 | 97.2 | 87.9 |
| CLIPFUSION (Ours) | 96.4 | 91.9 | 97.4 | 88.1 | |
| 4-shot | IIPAD | 97.0 | 91.2 | 97.4 | 88.3 |
| CLIPFUSION (Ours) | 96.8 | 92.5 | 98.1 | 88.6 | |
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPR | AUROC | AUPR | ||
| 1-shot | IIPAD | 94.2 | 97.2 | 85.4 | 87.5 |
| CLIPFUSION (Ours) | 94.5 | 97.3 | 85.5 | 88.1 | |
| 2-shot | IIPAD | 95.7 | 97.9 | 86.7 | 88.6 |
| CLIPFUSION (Ours) | 96.0 | 97.9 | 86.8 | 89.3 | |
| 4-shot | IIPAD | 96.1 | 98.1 | 88.3 | 89.6 |
| CLIPFUSION (Ours) | 96.4 | 98.2 | 88.3 | 90.0 | |
Appendix H Detailed quantitative results
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 99.20.0 100.00.0 97.40.7 99.40.4 98.20.9 100.00.0 98.51.0 99.20.3 99.30.3 100.00.0 98.80.2 99.20.3 99.30.4 100.00.0 Cable 86.50.0 87.30.0 57.74.6 88.84.2 88.91.9 95.00.9 62.35.9 91.02.7 88.40.7 96.30.4 70.06.1 91.02.7 90.90.9 96.70.4 Capsule 72.90.0 88.10.0 57.77.3 67.82.9 72.36.8 83.66.9 64.33.0 72.87.0 77.38.8 89.02.7 65.22.5 72.87.0 82.38.9 91.21.6 Carpet 100.00.0 99.90.0 96.61.0 95.30.8 99.80.3 100.00.0 97.80.5 96.60.5 99.80.3 100.00.1 97.90.4 96.60.5 100.00.0 100.00.0 Grid 98.80.0 99.90.0 54.26.7 63.610.3 99.50.3 99.40.4 67.24.2 67.78.3 99.40.2 99.60.4 68.13.8 67.78.3 99.60.1 99.50.4 Hazelnut 93.90.0 96.90.0 88.32.6 88.32.7 97.51.4 99.30.4 90.80.8 93.23.8 98.30.7 99.30.3 91.91.2 93.23.8 98.40.4 99.30.5 Leather 100.00.0 99.90.0 97.50.7 97.30.7 99.90.0 99.90.0 97.50.9 97.90.7 99.90.0 100.00.0 98.50.2 97.90.7 100.00.0 100.00.0 Metal nut 97.10.0 98.30.0 53.03.8 73.42.9 98.70.8 99.80.3 54.83.8 77.78.5 99.40.2 99.90.2 60.75.2 77.78.5 99.50.2 100.00.0 Pill 79.10.0 77.60.0 61.33.8 81.92.8 91.22.1 96.10.4 59.16.4 82.92.9 92.30.7 96.20.2 54.92.7 82.92.9 92.81.0 96.40.4 Screw 83.30.0 77.30.0 55.02.5 44.44.6 86.40.9 72.58.5 54.04.4 49.03.8 86.02.1 73.55.1 50.04.1 49.03.8 87.91.2 82.04.5 Tile 100.00.0 99.10.0 92.22.2 99.00.9 99.90.0 99.90.2 93.31.1 98.51.0 99.90.2 99.90.1 93.10.6 98.51.0 99.90.1 100.00.0 Toothbrush 87.50.0 96.90.0 82.51.2 83.33.8 92.24.9 98.91.6 87.64.2 85.93.5 97.51.6 99.10.8 89.22.5 85.93.5 96.72.6 98.50.8 Transistor 88.00.0 86.70.0 73.36.0 78.16.9 83.43.8 90.91.4 72.86.3 90.04.3 85.31.7 91.10.8 82.46.5 90.04.3 85.72.5 92.91.0 Wood 99.40.0 96.70.0 96.11.2 97.80.3 99.90.1 98.90.3 96.90.5 98.30.6 99.90.1 98.60.2 97.00.2 98.30.6 99.80.3 98.80.1 Zipper 91.50.0 94.00.0 85.82.7 92.30.5 88.85.9 96.50.7 86.32.6 94.02.1 94.01.4 95.32.2 88.32.0 94.02.1 94.50.5 96.20.6 Mean 91.80.0 93.20.0 76.63.1 83.43.0 93.12.0 95.40.8 78.93.1 86.33.3 94.41.3 95.90.3 80.42.5 88.82.6 95.21.3 96.80.2
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 99.80.0 100.00.0 99.20.2 99.80.1 99.40.3 100.00.0 99.60.3 99.80.1 99.80.1 100.00.0 99.70.0 99.80.1 99.80.1 100.00.0 Cable 91.20.0 92.70.0 64.93.8 93.82.2 93.21.1 97.10.5 69.66.6 95.11.3 92.90.6 97.80.3 76.15.6 97.10.7 94.40.3 98.10.2 Capsule 91.50.0 97.20.0 86.92.2 89.42.0 91.62.7 95.92.2 88.40.8 91.02.9 93.33.6 97.40.8 87.80.8 94.91.1 95.13.3 98.00.4 Carpet 100.00.0 100.00.0 99.00.2 98.70.2 99.90.1 100.00.0 99.40.1 99.00.1 99.90.1 100.00.0 99.40.1 98.80.2 100.00.0 100.00.0 Grid 99.60.0 100.00.0 75.03.3 81.14.9 99.90.1 99.80.1 82.52.3 84.14.0 99.80.1 99.90.1 83.01.8 86.44.0 99.90.0 99.80.1 Hazelnut 96.90.0 98.30.0 93.31.7 92.92.2 98.60.7 99.60.2 94.10.5 96.02.0 99.10.4 99.60.1 94.80.6 97.01.2 99.10.2 99.60.3 Leather 100.00.0 100.00.0 99.20.2 99.10.2 100.00.0 100.00.0 99.20.3 99.30.2 100.00.0 100.00.0 99.60.1 99.60.1 100.00.0 100.00.0 Metal nut 99.30.0 99.60.0 82.02.7 91.01.1 99.70.2 99.90.1 82.21.4 92.34.0 99.90.0 100.00.0 85.51.7 97.02.6 99.90.1 100.00.0 Pill 95.70.0 94.80.0 88.31.3 96.50.6 98.30.5 99.30.1 87.92.6 96.60.7 98.60.1 99.30.0 87.01.2 96.90.4 98.60.2 99.40.1 Screw 93.10.0 89.30.0 78.11.0 71.42.3 94.20.6 87.54.7 77.31.3 72.93.4 94.11.5 87.43.2 75.72.8 71.81.9 94.90.8 93.02.3 Tile 100.00.0 99.60.0 97.20.7 99.60.3 100.00.0 100.00.1 97.60.4 99.40.4 100.00.1 100.00.0 97.60.2 99.60.1 100.00.0 100.00.0 Toothbrush 95.60.0 98.80.0 93.70.5 93.51.4 96.72.0 99.60.5 95.21.6 94.11.4 99.00.6 99.70.3 95.80.7 94.80.7 98.71.1 99.50.2 Transistor 87.10.0 84.70.0 66.27.5 77.75.5 79.04.0 89.81.5 69.06.5 89.33.9 80.72.3 89.90.7 77.68.4 84.59.0 80.73.2 92.31.3 Wood 99.80.0 99.00.0 98.80.3 99.30.1 100.00.0 99.70.1 99.00.1 99.50.2 100.00.0 99.60.1 99.10.0 99.50.2 99.90.1 99.70.0 Zipper 97.50.0 97.90.0 95.50.9 97.20.3 96.81.8 98.90.3 95.41.0 97.81.0 98.30.4 98.31.0 96.20.8 99.10.7 98.50.2 98.70.2 Mean 96.50.0 96.80.0 88.11.7 92.21.5 96.50.9 97.80.4 89.31.7 93.81.7 97.00.7 97.90.2 90.51.6 94.51.5 97.30.6 98.50.1
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 89.50.0 94.10.0 96.10.5 97.90.1 97.50.2 97.30.2 96.90.1 98.10.0 97.70.1 97.50.0 97.10.1 98.20.0 97.80.0 97.70.1 Cable 77.00.0 73.20.0 88.41.2 95.50.8 93.80.6 91.30.2 90.00.8 96.40.3 94.30.4 92.20.3 92.10.4 97.50.3 94.90.1 92.90.4 Capsule 86.90.0 94.60.0 94.50.6 95.60.4 94.60.8 97.90.4 95.20.5 96.50.4 96.40.3 98.40.2 96.20.4 96.80.6 96.20.5 98.50.1 Carpet 95.40.0 99.30.0 97.80.2 98.40.1 99.40.0 99.40.1 98.20.0 98.50.1 99.30.0 99.50.0 98.40.0 98.60.1 99.30.0 99.50.0 Grid 82.20.0 98.90.0 70.22.8 58.84.9 96.81.0 98.80.2 70.82.0 62.63.2 97.70.8 98.90.4 77.01.8 69.41.3 98.00.2 99.20.1 Hazelnut 94.30.0 96.70.0 95.40.7 95.80.6 98.50.2 98.20.1 96.80.3 96.30.6 98.70.1 98.20.1 97.20.2 97.60.1 98.80.0 98.70.1 Leather 96.70.0 99.20.0 98.50.1 98.80.2 99.30.0 99.30.0 98.70.1 99.00.1 99.30.0 99.30.0 98.80.0 99.10.0 99.30.0 99.40.0 Metal nut 61.00.0 74.20.0 74.61.1 89.31.4 90.00.6 86.21.2 80.32.1 94.61.4 91.40.4 89.70.9 82.73.9 95.91.8 92.90.4 91.30.8 Pill 80.00.0 83.00.0 84.81.0 93.11.1 96.40.3 95.50.2 87.30.7 94.20.3 97.00.2 95.60.2 88.90.5 94.80.4 97.10.0 95.80.2 Screw 89.60.0 97.20.0 83.30.7 89.60.5 94.50.4 95.60.4 89.80.8 90.00.7 95.20.3 95.60.6 90.80.2 91.31.0 96.00.5 96.30.5 Tile 77.60.0 96.20.0 84.11.1 94.10.5 96.30.2 97.70.1 87.70.2 94.40.2 96.50.1 97.80.1 88.90.3 94.60.1 96.60.1 97.80.1 Toothbrush 86.90.0 93.20.0 97.30.3 97.30.4 97.80.1 99.10.1 97.70.3 97.50.2 98.10.1 99.20.1 98.40.2 98.40.4 98.40.5 99.20.1 Transistor 74.70.0 69.20.0 90.22.8 84.92.7 85.01.8 87.32.6 92.32.1 89.60.9 88.31.0 88.60.8 94.02.7 90.71.4 88.51.2 90.30.6 Wood 93.40.0 96.00.0 90.70.4 92.70.9 94.61.0 95.71.2 91.90.1 93.20.7 95.30.4 96.40.3 92.20.1 93.50.3 95.40.2 96.40.4 Zipper 91.60.0 98.20.0 93.90.8 97.40.4 93.90.8 98.40.3 95.40.3 98.00.1 94.10.7 98.30.1 96.10.2 98.10.1 94.20.4 98.50.1 Mean 85.10.0 90.90.0 89.30.9 92.01.0 95.20.5 95.80.1 91.30.7 93.30.6 96.00.3 96.30.1 92.60.7 94.30.5 96.20.3 96.80.1
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 76.40.0 88.40.0 89.80.8 93.50.3 91.20.4 93.70.4 91.70.2 93.90.3 91.80.3 94.10.1 92.20.2 94.00.2 91.60.2 94.50.3 Cable 42.90.0 71.50.0 59.13.2 84.71.0 72.52.3 79.70.8 66.52.8 88.50.9 74.72.3 82.01.0 74.21.8 91.70.6 77.01.1 84.41.0 Capsule 62.10.0 88.50.0 80.02.0 83.90.9 85.62.7 92.51.1 82.32.1 86.61.0 90.60.6 94.10.2 85.71.3 87.81.9 90.11.5 94.70.5 Carpet 84.10.0 98.20.0 92.90.3 93.30.3 97.40.4 98.30.1 93.90.2 93.70.4 97.30.3 98.40.1 94.40.2 93.90.4 97.00.2 98.50.1 Grid 57.00.0 95.70.0 41.24.6 21.79.5 90.52.7 96.11.1 45.13.6 23.73.8 92.82.5 97.10.9 55.53.4 30.44.6 93.60.6 96.90.7 Hazelnut 81.60.0 93.80.0 85.71.9 88.31.3 93.70.9 94.30.3 89.40.9 89.81.3 94.20.3 94.70.5 90.40.7 92.00.3 94.20.3 95.60.6 Leather 91.10.0 97.70.0 95.60.2 95.21.0 98.60.0 97.80.1 96.20.2 95.90.3 98.30.4 97.70.1 96.30.1 96.40.1 98.00.4 97.60.1 Metal nut 31.80.0 74.10.0 38.11.6 66.72.9 84.71.1 80.11.6 48.25.0 79.64.2 86.70.8 85.81.6 54.08.8 83.85.5 89.40.1 87.61.1 Pill 65.00.0 87.60.0 78.90.6 89.51.6 93.50.2 93.90.3 84.30.4 91.60.5 94.50.2 94.20.2 86.60.4 92.50.4 94.60.3 94.50.1 Screw 68.50.0 89.50.0 51.61.7 68.11.3 82.31.1 84.00.8 69.52.1 69.02.1 84.10.5 83.61.5 72.30.8 72.43.1 86.31.8 85.91.1 Tile 51.20.0 91.90.0 66.71.5 82.51.1 89.40.4 94.50.2 71.90.5 82.50.5 89.60.4 94.70.2 73.60.9 83.00.1 89.90.3 94.80.1 Toothbrush 67.70.0 90.60.0 82.11.5 79.02.4 85.31.0 91.60.9 83.32.6 81.00.7 84.71.4 92.10.3 87.11.7 85.53.0 86.03.3 92.70.4 Transistor 43.40.0 52.50.0 70.37.0 70.94.6 65.01.8 68.83.0 76.55.5 78.81.5 68.61.1 71.30.6 82.27.4 79.52.8 69.01.1 73.11.3 Wood 74.10.0 93.10.0 86.50.6 87.11.0 91.00.6 95.00.8 88.00.2 86.81.4 91.80.6 95.30.2 88.40.2 87.70.4 91.70.3 95.30.4 Zipper 71.70.0 94.00.0 81.72.0 91.21.1 86.01.7 95.40.6 85.60.7 92.80.4 86.41.6 95.70.3 87.20.8 93.40.2 86.90.7 96.00.3 Mean 64.60.0 87.10.0 73.32.0 79.72.0 87.11.2 90.40.1 78.21.8 82.31.3 88.40.9 91.40.2 81.31.9 84.31.6 89.00.8 92.10.2
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Candle 95.40.0 94.30.0 70.84.1 85.11.4 93.41.4 92.60.7 75.82.1 85.31.5 94.81.0 93.41.8 77.51.6 87.80.8 95.10.3 94.00.9 Capsules 85.00.0 80.30.0 51.07.8 60.07.6 85.03.1 83.01.2 51.74.6 57.85.4 84.90.8 83.10.9 52.73.4 63.45.4 86.81.7 83.00.8 Cashew 92.10.0 81.70.0 62.39.9 89.54.4 94.00.4 89.41.7 74.63.6 93.60.6 94.30.5 89.31.0 77.73.2 93.01.5 95.20.8 90.81.4 Chewinggum 96.50.0 96.80.0 69.94.9 97.30.3 97.60.8 97.40.3 82.72.1 97.80.6 97.30.8 97.60.3 83.53.7 98.30.3 97.70.3 97.70.3 Fryum 80.30.0 88.40.0 58.35.9 75.04.8 88.51.9 92.21.0 69.29.0 83.42.4 90.50.4 93.11.2 71.25.9 88.61.3 90.80.5 93.70.4 Macaroni1 76.20.0 79.30.0 62.14.6 68.03.4 82.91.5 86.80.7 62.25.0 75.64.6 83.31.9 87.91.1 65.93.9 82.92.7 85.20.9 88.91.4 Macaroni2 63.70.0 61.80.0 47.55.9 55.64.6 70.20.9 72.22.1 50.82.9 57.35.6 71.82.0 70.82.2 55.02.9 61.71.8 70.92.2 72.71.8 PCB1 73.60.0 83.70.0 76.21.2 78.91.1 75.623.0 91.31.3 62.410.8 71.520.0 76.75.2 89.74.0 82.61.5 84.76.7 88.31.7 89.72.8 PCB2 51.20.0 61.60.0 61.22.0 81.50.8 62.23.9 69.75.3 66.82.0 84.31.7 62.63.7 75.93.3 73.52.4 84.31.0 67.52.6 76.41.4 PCB3 73.40.0 67.80.0 51.412.2 82.72.3 74.11.1 76.64.5 67.33.8 84.81.2 78.81.9 79.22.7 65.91.9 87.01.1 83.31.7 78.41.2 PCB4 79.60.0 76.90.0 76.13.6 93.92.8 85.28.9 91.41.9 69.313.7 94.33.2 82.39.9 87.85.0 85.42.0 95.61.6 87.68.0 94.11.6 Pipe fryum 69.70.0 81.50.0 66.72.2 90.71.7 97.21.1 95.81.1 75.31.8 93.51.3 98.00.6 96.60.6 82.92.2 96.40.7 98.50.4 97.00.2 Mean 78.10.0 79.50.0 62.85.4 79.92.9 83.84.0 86.50.1 67.45.1 81.64.0 84.62.4 87.00.5 72.82.9 85.32.1 87.31.8 88.00.2
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Candle 95.80.0 94.90.0 69.23.9 86.62.3 93.61.5 92.90.6 72.81.0 86.81.7 95.11.1 93.91.8 72.51.1 88.91.1 95.30.4 94.41.0 Capsules 90.90.0 88.80.0 63.45.7 72.35.3 89.92.5 90.40.7 63.42.0 73.64.7 88.90.7 90.60.4 63.02.3 78.43.1 91.51.4 90.70.3 Cashew 96.40.0 91.90.0 78.25.7 94.62.0 97.20.2 95.30.8 86.12.2 96.90.3 97.30.2 95.20.4 88.42.0 96.50.7 97.70.4 95.80.6 Chewinggum 98.60.0 98.70.0 79.83.6 98.90.1 99.00.3 98.90.1 89.51.9 99.10.2 98.90.3 99.00.1 88.53.2 99.30.1 99.00.1 99.00.1 Fryum 90.10.0 94.60.0 74.52.9 87.62.4 94.71.0 96.60.4 81.05.4 92.11.3 95.80.2 96.90.5 81.53.0 95.00.6 96.00.3 97.10.1 Macaroni1 75.80.0 81.10.0 60.42.9 67.83.4 84.91.2 88.40.9 63.14.3 74.95.2 84.71.5 88.90.9 64.92.1 82.13.5 86.50.6 89.81.3 Macaroni2 60.30.0 59.50.0 51.75.0 54.93.2 68.41.8 71.02.4 52.71.5 57.22.6 70.41.8 68.34.9 54.92.5 60.23.0 69.62.8 70.93.3 PCB1 78.40.0 87.00.0 68.62.4 72.12.5 76.519.0 91.21.4 60.47.7 72.616.4 78.34.3 89.24.2 77.42.9 81.09.2 87.71.7 89.33.5 PCB2 49.20.0 66.80.0 63.31.2 84.40.4 64.93.3 69.85.1 68.92.6 86.61.1 65.84.0 75.73.2 75.01.7 86.21.0 71.33.4 75.52.0 PCB3 76.50.0 70.30.0 52.310.8 84.61.5 73.51.6 79.53.3 65.23.8 86.10.5 80.91.6 80.92.9 64.52.4 88.31.1 84.81.8 81.01.4 PCB4 77.70.0 78.50.0 74.72.6 92.83.1 78.515.5 88.83.1 67.611.9 93.23.4 72.516.2 84.85.6 84.02.0 94.91.2 85.68.9 91.92.1 Pipe fryum 82.30.0 90.50.0 79.21.5 95.40.6 98.60.5 98.00.6 84.51.7 96.80.7 99.00.3 98.40.3 89.81.7 98.30.3 99.20.2 98.60.1 Mean 81.20.0 83.60.0 68.34.0 82.82.3 85.14.0 88.40.2 71.63.8 84.83.2 85.82.7 88.50.5 75.62.2 87.52.1 88.81.8 89.50.1
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Candle 88.90.0 97.50.0 91.72.2 97.20.2 97.40.2 98.40.2 94.90.8 97.70.3 97.70.1 98.50.1 95.40.2 97.90.1 97.80.2 98.60.1 Capsules 81.60.0 96.10.0 70.91.1 93.20.9 96.40.6 97.80.3 75.71.7 94.00.2 96.80.3 97.90.5 79.10.7 94.80.5 97.10.2 97.90.1 Cashew 84.70.0 72.50.0 95.50.6 98.10.1 98.50.2 98.20.3 96.40.4 98.20.2 98.50.1 98.50.1 97.20.3 98.30.2 98.70.0 98.00.5 Chewinggum 93.30.0 99.10.0 90.10.4 96.90.3 98.60.1 99.00.1 93.10.7 96.60.1 98.60.1 98.90.1 94.40.5 96.80.1 98.50.1 98.80.1 Fryum 88.50.0 92.50.0 93.30.6 93.30.5 96.40.3 95.20.3 94.10.6 94.00.3 97.00.2 96.10.3 95.00.4 94.20.2 97.10.1 96.50.4 Macaroni1 70.90.0 98.60.0 89.40.9 95.20.4 96.40.6 99.10.1 91.70.3 96.01.3 96.50.7 99.20.1 93.50.5 97.00.3 97.00.2 99.30.1 Macaroni2 59.30.0 96.50.0 86.41.1 89.11.6 96.80.4 97.40.3 90.10.8 90.21.9 96.80.6 97.50.1 90.20.3 93.90.3 97.30.3 97.90.2 PCB1 61.20.0 91.60.0 89.90.3 96.11.5 96.60.6 95.50.2 90.60.6 97.60.9 97.00.9 96.20.2 93.21.5 98.11.0 98.10.9 97.40.9 PCB2 71.60.0 91.10.0 90.91.4 95.40.2 93.00.4 96.10.3 93.90.9 96.00.3 93.90.2 96.70.1 93.71.0 96.60.2 94.60.4 97.30.3 PCB3 85.30.0 88.40.0 93.90.3 96.20.3 94.30.3 96.20.2 95.10.5 97.10.1 95.10.2 96.90.2 95.70.1 97.40.2 95.80.1 97.30.2 PCB4 94.40.0 93.50.0 89.60.6 95.60.6 94.00.9 95.80.5 90.70.9 96.20.4 95.60.3 96.70.2 92.10.5 97.00.2 96.10.3 97.30.2 Pipe fryum 75.40.0 87.60.0 97.20.6 98.80.2 98.30.2 99.40.1 98.10.4 99.10.1 98.50.2 99.40.1 98.50.1 99.10.0 98.70.1 99.50.0 Mean 79.60.0 92.10.0 89.90.8 95.40.6 96.40.4 97.30.1 92.00.7 96.10.5 96.80.3 97.70.1 93.20.5 96.80.3 97.20.2 98.00.1
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Candle 83.50.0 93.30.0 81.55.3 92.60.4 94.00.4 94.50.7 87.31.2 93.40.6 94.20.2 94.80.3 88.30.7 94.10.4 94.40.2 95.30.3 Capsules 35.30.0 83.70.0 30.61.1 66.64.5 73.63.5 87.73.1 38.43.7 67.92.3 75.91.9 88.02.6 43.32.0 69.03.2 77.01.4 87.20.7 Cashew 76.40.0 92.30.0 73.42.1 90.80.2 91.10.8 91.71.4 78.42.7 91.41.0 90.40.6 90.11.2 81.22.8 92.10.3 91.30.9 89.61.5 Chewinggum 70.40.0 89.10.0 58.10.6 78.21.3 91.00.5 85.91.2 63.72.4 78.00.4 90.90.7 85.51.0 67.21.8 79.30.8 91.00.4 84.50.7 Fryum 77.40.0 88.30.0 71.11.6 78.72.3 89.11.0 83.22.1 71.20.8 81.42.8 89.30.2 84.11.7 73.21.3 81.01.2 89.70.5 85.31.4 Macaroni1 34.30.0 87.70.0 62.24.4 83.41.3 84.62.3 90.00.5 71.82.4 86.24.6 85.21.4 91.20.8 76.62.1 89.60.7 86.80.8 91.80.7 Macaroni2 21.40.0 75.50.0 54.93.6 66.03.0 89.32.4 80.32.0 65.63.4 67.26.5 88.61.7 79.21.5 65.91.5 78.30.9 90.51.3 83.01.8 PCB1 26.30.0 73.50.0 63.91.8 79.010.7 82.56.0 86.30.7 68.44.1 86.11.7 83.85.0 86.12.9 70.23.3 88.12.6 87.92.1 84.92.8 PCB2 37.20.0 73.50.0 64.43.8 80.90.5 73.61.5 80.51.8 72.93.4 82.91.8 76.20.9 82.80.9 71.92.6 83.71.0 78.01.3 84.01.9 PCB3 56.10.0 66.20.0 69.01.2 78.12.0 79.52.5 82.71.9 74.02.3 82.21.1 82.31.8 83.11.6 77.20.8 84.41.9 84.21.0 81.73.1 PCB4 80.40.0 81.20.0 59.11.8 77.93.1 76.64.1 84.51.8 62.63.6 79.54.8 81.71.2 86.11.2 67.92.6 83.52.5 84.20.7 88.51.0 Pipe fryum 82.30.0 95.20.0 83.90.8 93.60.5 96.10.6 96.60.3 86.90.9 94.50.4 96.20.6 97.20.2 88.71.3 95.00.5 96.60.2 97.00.2 Mean 56.80.0 83.30.0 64.32.4 80.52.5 85.12.1 87.00.3 70.12.6 82.62.3 86.21.4 87.30.2 72.61.9 84.91.4 87.60.9 87.70.5
Appendix I Detailed qualitative results